Laser light of wavelength 632.8 nm falls normally on a slit that is 0.0240 mm wide. The transmitted light is viewed on a distI need B and C
Answer #1

Solution:

Given data

The wavelength of the incident light, X= 632.8 nm

Width of the slit, d=0.0240\; \mathrm{mm}

The intensity at the center of the central bright fringe, 1. = 8.90 W/m

(b)

The maximum order of diffraction can be calculated as follows

sin ma d

\frac{m\lambda }{d}=1\; \; \; \; \; \; \; \; \; (\sin \theta =1)

m=\frac{d}{\lambda }

m=\frac{0.024\times 10^{-3}}{632.8\times 10^{9} }

m=37.9\approx 37

Now, we calculate the angle corresponding to this order as follows

sin ma d

\theta =\sin ^{-1}\left ( \frac{m\lambda }{d} \right )

\theta_{37} =\sin ^{-1}\left ( \frac{37\times 632.8\times 10^{-9}}{0.024\times 10^{-3}} \right )

\theta_{37} =77.42^{\circ}

Now, calculate the angle corresponding to m = 36

\theta_{36} =\sin ^{-1}\left ( \frac{36\times 632.8\times 10^{-9}}{0.024\times 10^{-3}} \right )

\theta_{36} =71.62^{\circ}

Center of the fringe can be calculated as follows

\theta_{max} =\frac{77.42+71.62}{2}

\theta_{max} =74.52^{\circ}

(c)

The intensity of the bright fringe before the dark fringe has seen in part (b) can be calculated as follows

\beta =\frac{2\pi }{\lambda }d\sin \theta

\beta =\frac{2\times 3.14}{632.8\times 10^{-9} }(0.024\times 10^{-3})\sin 74.52

\beta =229

Now, the intensity of the bright fringe can be calculated as follows

I=I_{o}\frac{(\sin \frac{\beta }{2})^{2}}{(\frac{\beta }{2})^{2}}

I=I_{o}\frac{(\sin \frac{229 }{2})^{2}}{(\frac{229 }{2})^{2}}

I=I_{o}\times 6.32\times 10^{-5}

We know,

1. = 8.90 W/m

Then we have

I=8.90\times 6.32\times 10^{-5}

I=56.25\times 10^{-5}

I=5.625\times 10^{-4}\; \mathrm{W/m^{2}}